Run Planning Meeting Notes, Apr12-Apr 18, 2018

From GlueXWiki
Jump to: navigation, search

<- Back to Run Coordination Meetings:Fall2017 Spring2018 Run

Thursday, April 12 at 08:45

  1. Notes from the previous 24 hours:
    • TAC run (evening strip chart) was finished around 10pm. Took data with both the 750um and 75um converter.
    • Production started around 11 pm and has continued through the night.
  2. Issues:
    • When adjusting a procedure, please contact the expert before doing so to make sure the changes you make are okay.
    • ~1hr 20 mins of DAQ issues while the target was filling and we could not run in raw mode.
    • PS scalers (Alexandre)
  1. Reports:
  2. Run Plan
    • Pass change this morning
      • Short access to install copper and tungsten shields
      • Move muon chamber back into beam line
    • Production running
      • Harp scan when beam comes back
    • Fast raster scan postponed to Monday
    • Any more systematic studies we want to do?
  3. AOB

Friday, April 13 at 08:45

  1. Notes from the previous 24 hours:
  2. Issues:
    • Bill noticed some minor inconsistencies with monitoring plots (beta vs. p, beam energy, and trigger bit)
  3. Reports:
  4. Run Plan
  1. AOB
    • Have Lubomir for opportunistic access to remove Cu and half of W shield- any others? Beam studies on Monday and RF recovery Tuesday
    • Run meeting this weekend?

Saturday, April 14 at 08:45

No meeting

Sunday, April 15 at 08:45

No meeting

Monday, April 16 at 08:45

  1. Notes from the previous 48 hours:
  1. Issues:
  2. Reports
  3. Run Plan
    • 8hr beam studies today
      • Controlled access to adjust shielding, replace AC voltage adapter, FCAL bases, and other technician work
    • Finish materials scan
    • Production this evening
    • RF recovery tomorrow
      • Switch AC to 3.5 mm collimator
        • Take data through Thurs. with this setting: 150 nA, 75 micron converter, 0/90, AMO
        • proper beam tuning?
    • Best time to try to try the low current lock on the AC for the next TAC run? Tentatively postponed until Wed. before raster tests
  1. AOB

Tuesday, April 17 at 08:45

  1. Notes from the previous 48 hours:
  1. Issues:
  2. Reports
  1. Run Plan
    • RF Recovery
      • Switch AC to 3.5 mm collimator
        • Settings to run at from beam line group? Last I heard was 75 um converter, ~300nA (normalize DAQ rates), 0/90 PARA/PERP x2, AMO x2
        • Talked to Todd about a performing a beam convergence. Will be available all day tomorrow
      • Reinstall shielding
  1. AOB

Wednesday, April 18 at 08:45

  1. Notes from the previous 48 hours:
    • Moved AC to 3.5 mm collimator
    • RF recovery took longer than expected
    • Last night once we got beam back, it was seen that we could not lock on the AC. We went into controlled access and have a radcon escort into the collimator cave so Alex B. could fix AC. He first thought it was the traditional voltage adapter problem, but it turned out to be a cable issue.
  2. Issues:
    • Goniometer unresponsive- vertical stage trip
    • Cable issue on AC (Alex B.)
    • Another issue with the AC lock: Notice that the Active Collimator (AC) lock is not on. MCC tells us that the AC is showing an invalid status and cannot be used for beam locking.
      • Concern that the optimal AC position lock for the 3.4 mm collimator may not be the same as that for the 5 mm collimator.
      • Beam position was checked with the gamma profiler, even though there is no radiator plate attached. Rates look OK.
      • After tests look OK, the crew decided to go ahead with the program, trusting that the AC lock is roughly centers the beam on the 3.4 mm collimator.
    • Coffee maker broken (Mark I.)
    • Noisy CDC channel
  3. Reports
  4. Run Plan
    • Running with thin diamond
      • With Sasha on shift last night, it was decided to run by normalizing to PS rates
      • Do not run above 500nA until some counters are turned off
    • What else would we like to do with the thin diamond?
    • Beam tune, fast raster scan, and low current AC lock test?
  5. AOB
    • Mark D. takes over as RC